Question1

The conditions in a certain electrical circuit is represented by the following differential equation:

1/50Q’’ +Q’ +8q =50 Cos30t

1.      Determine an expression for q in term of t

2.      Determine an expression for the current  I        (hint I =dq/dt)

3.      Determine the amplitude and the frequency of the steady state current.

 

Question2

Find the general solution of the following DE using the method of undetermined coefficients :

Y’’ +2y’+5y = 34sinxcosx  


1
Expert's answer
2021-07-15T09:36:51-0400

Question1

1) "\frac{1}{{50}}q'' + q' + 8q = 50\cos 30t"

"q'' + 50q' + 400q = 2500\cos 30t"

characteristic equation:

"{k^2} + 50k + 400 = 0"

"D = 2500 - 1600 = 900"

"{k_1} = \frac{{ - 50 - 30}}{2} = - 40"

"{k_2} = \frac{{ - 50 + 30}}{2} = - 10"

Then the general solution of the homogeneous equation is

"{q_0} = {C_1}{e^{ - 40t}} + {C_2}{e^{ - 10t}}"

We will seek a particular solution in the form

"\widetilde q = A\cos 30t + B\sin 30t \Rightarrow {\widetilde q^\prime } = - 30A\sin 30t + 30B\cos 30t \Rightarrow {\widetilde q^{\prime \prime }} = \\=- 900A\cos 30t - 900B\sin 30t"

Substitute the obtained values ​​into the original equation:

"- 900A\cos 30t - 900B\sin 30t - 1500A\sin 30t + 1500B\cos 30t + 400A\cos 30t + 400B\sin 30t = 2500\cos 30t"

"( - 900A + 1500B + 400A)\cos 30t + ( - 900B - 1500A + 400B)\sin 30t = 2500\cos 30t"

"\left\{ \begin{array}{l}
 - 500A + 1500B = 2500\\
 - 1500A - 500B = 0
\end{array} \right. \Rightarrow"

"\Rightarrow \left\{ {\begin{matrix}
{A = - \frac{1}{2}}\\
{B = \frac{3}{2}}
\end{matrix}} \right."

So,

"\widetilde q = - \frac{1}{2}\cos 30t + \frac{3}{2}\sin 30t"

"q(t) = {q_0} + \widetilde q = {C_1}{e^{ - 40t}} + {C_2}{e^{ - 10t}} - \frac{1}{2}\cos 30t + \frac{3}{2}\sin 30t"

Answer: "q(t) = {C_1}{e^{ - 40t}} + {C_2}{e^{ - 10t}} - \frac{1}{2}\cos 30t + \frac{3}{2}\sin 30t"

2)

"I(t) = \frac{{dq}}{{dt}} = {\left( {{C_1}{e^{ - 40t}} + {C_2}{e^{ - 10t}} - \frac{1}{2}\cos 30t + \frac{3}{2}\sin 30t} \right)^\prime } =\\= - 40{C_1}{e^{ - 40t}} - 10{C_2}{e^{ - 10t}} + 15\sin 30t + 45\cos 30t"

Answer: "I(t) = - 40{C_1}{e^{ - 40t}} - 10{C_2}{e^{ - 10t}} + 15\sin 30t + 45\cos 30t"

3) The steady state solution is "\widetilde q = - \frac{1}{2}\cos 30t + \frac{3}{2}\sin 30t = \sqrt {\frac{5}{2}} \sin \left( {30t - arctan\frac{1}{3}} \right)"

So, the amplitude is "A = \sqrt {\frac{5}{2}}" , the frequency is "f =\frac{ 30}{2\pi}"

Answer: "A = \sqrt {\frac{5}{2}}" , "f = \frac{30}{2 \pi}"

Question 2

"y'' + 2y' + 5y = 34\sin x\cos x"

"y'' + 2y' + 5y = 17\sin 2x"

characteristic equation:

"{k^2} + 2k + 5 = 0"

"D = 4 - 20 = - 16"

"{k_1} = \frac{{ - 2 - 4i}}{2} = - 1 - 2i"

"{k_2} = \frac{{ - 2 + 4i}}{2} = - 1 + 2i"

Then the general solution of the homogeneous equation is

"{y_0} = {e^{ - x}}\left( {{C_1}\cos 2x + {C_2}\sin 2x} \right)"

We will seek a particular solution in the form

"\widetilde y = A\sin 2x + B\cos 2x \Rightarrow {\widetilde y^\prime } = 2A\cos 2x - 2B\sin 2x \Rightarrow {\widetilde y^{\prime \prime }} = - 4A\sin 2x - 4B\cos 2x"

Substitute the obtained values ​​into the original equation:

"- 4A\sin 2x - 4B\cos 2x + 4A\cos 2x - 4B\sin 2x + 5A\sin 2x + 5B\cos 2x = 17\sin 2x"

"(A - 4B)\sin 2x + (4A + B)\cos 2x = 17\sin 2x"

"\left\{ \begin{array}{l}
A - 4B = 17\\
4A + B = 0
\end{array} \right. \Rightarrow A = 1,B = - 4"

So,

"\widetilde y = \sin 2x - 4\cos 2x"

"y = {y_0} + \widetilde y = {e^{ - x}}\left( {{C_1}\cos 2x + {C_2}\sin 2x} \right) + \sin 2x - 4\cos 2x"

Answer: "y = {e^{ - x}}\left( {{C_1}\cos 2x + {C_2}\sin 2x} \right) + \sin 2x - 4\cos 2x"